OpenStudy (anonymous):

A collection of quarters and nickels is worth $13.70. There are 106 coins in all. How many of each coin are there? quarters nickels

OpenStudy (anonymous):

are you in k12

OpenStudy (danjs):

same type question... you got this!

OpenStudy (danjs):

type what you have so far

OpenStudy (anonymous):

yes I am I think 64 nickels and 42 quarters

OpenStudy (anonymous):

@1497 64 nickels 42 quarters

OpenStudy (anonymous):

yup your right

OpenStudy (danjs):

you setting up 2 equations?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{q+n=106\\ 5n+25q=13.70\times 100}\end{align}\)

OpenStudy (mathmath333):

convert it into cents

OpenStudy (danjs):

right, i was wondering if @1497 could do that, since this is lie the 5th problem of this type.